A​ "normal" temperature for a certain animal is 94.8
degrees.If a temperature x that differs from the normal by at least 2.4 degrees
is considered​ unhealthy, write the condition for an unhealthy temperature x as an inequality involving an absolute​ value, and solve for x.Write the condition for an unhealthy temperature x as an inequality involving an absolute value.
The absolute value is |x-94.8|>=2.4 I just don't know what to do after this to get the solution.

Answer by ikleyn(52832)

.
|x - 94.8| >= 2.4.


The solutions for this inequality are those values of x that are remoted farther than 2.4 from the point (from the value) of 94.8.


So, these solutions are these two sets  

    x >= 94.8 + 2.4 = 97.2

and

    x <= 94.8 - 2.4 = 92.4.


Answer.  The solutions to the given inequality are these two sets  {x <= 92.4}  and  {x >= 97.2}.


         In other words, the solution is the union of these two sets.

Is everything clear to you now ?